### v2

# Novel thermodynamic inequality for rotating AdS black holes

Abstract
We propose a new thermodynamic inequality for stationary and asymptotically Anti-de Sitter rotating black holes, $4\pi J^{2}/(3MV)<1$. This inequality is derived through an analysis of the roots of the identity which holds among the thermodynamic variables to avoid a naked singularity. We have analyzed the Kerr-AdS black hole and the asymptotically AdS uncharged and rotating black strings and find strong supporting evidence for different horizon topologies. Using this inequality we verify the validity of reverse isoperimetric inequality in its conventional form (${\cal R}\geq 1$) in the presence of rotation. By examining a wide range of black hole solutions, we confirm the validity of the proposed inequality. Assuming that reverse isoperimetric inequality continues to hold in higher dimensions, and employing the intermediate refined reverse isoperimetric inequalities introduced in [Phys. Rev. Lett. 131, 241401] as guiding principles, we extend the inequality to higher-dimensional spacetimes.
Introduction. The thermodynamic properties of black holes have played an outstanding role in advancing our understanding of the interplay between gravitation, quantum theory, and thermodynamics because of the seminal works of Bekenstein and Hawking [1, 2]. Bekenstein first proposed that the entropy of a black hole is proportional to the area of its event horizon [1], a notion that was profoundly supported by Hawking’s discovery of black hole radiation [2]. In Anti-de Sitter (AdS) spacetimes, these ideas acquire additional depth due to the AdS/CFT correspondence [3], which facilitates the study of gravitational systems through dual conformal field theories, thereby offering new perspectives on entropy bounds.
Recent developments have significantly expanded the classical picture of black hole thermodynamics by incorporating notions such as thermodynamic volume [4, 5], critical behavior in extended phase space [6], and the intriguing concept of black holes functioning as heat engines [7]. This broader framework, often referred to as “black hole chemistry” [8], has motivated investigations into entropy inequalities and phase structures within the AdS/CFT correspondence. Additionally, recent work on holographic complexity [9], the role of chemical potentials in holographic thermodynamics [10], and the higher-dimensional origins of black hole thermodynamics [11, 12] has further highlighted the necessity for rigorous mathematical treatments of entropy bounds in AdS black hole spacetimes. Initial efforts to explore these ideas in rotating black holes and black rings demonstrated the existence of nontrivial phase behavior and thermodynamic volumes in complex horizon geometries [13]. Further studies extended these ideas to de Sitter backgrounds [14], accelerating spacetimes [15], and scenarios involving holography and information geometry. While some literatures include possible violations of the RII, a conclusive counterexample to the conjecture is still lacking.
A central concept arising from this framework is the reverse isoperimetric inequality, which conjectures that for a given thermodynamic volume, there exists a lower bound on the black hole entropy [14, 16]. While this inequality holds for many conventional black hole solutions, a number of important exceptions have been identified, particularly in the context of ultraspinning and super-entropic black holes [17, 20]. These solutions violate the inequality, suggesting that entropy may exceed the expected geometric bounds, thus challenging our intuition about gravitational thermodynamics and entropy-area relations. Such violations have motivated inquiries into the underlying conditions that lead to such behavior, as well as into the stability and physical realizability of these exotic solutions [17, 23].
Simultaneously, efforts have been made to develop a microscopic and holographic interpretation of thermodynamic volume and its relation to the degrees of freedom of the dual theory [18]. These developments raise important questions about the fundamental nature of volume, entropy, and temperature in gravitational theories, particularly in lower-dimensional settings [25] and in theories with higher-curvature corrections or Chern-Simons terms [24]. Moreover, investigations into the geometry of black hole horizons with unusual topologies [19], and the structure of gauged supergravity solutions [22], further illustrate the diversity and complexity of black hole chemistry. Refined versions of the quantum Penrose and reverse isoperimetric inequalities, valid for all known three-dimensional asymptotically AdS quantum black holes are also proposed in [26]. This paper aims to prpose thermodynamic inequality for uncharged, stationary and asymptotically Anti-de Sitter rotating black holes.
Evidence for the inequality. Consider first the Kerr-AdS black hole [27, 30, 31], for which the thermodynamics in extended phase space was first studied in [5]. For this case, the following identity holds among the thermodynamic parameters [32] (see Eq. (37) of Supplement)
$36\pi M^{2}V^{2}-M^{2}A^{3}-64\pi^{3}J^{4}=0,$ (1)
which can be solved explicitly for the horizon area as
$A^{3}=36\pi V^{2}\left[1-\left(\frac{4\pi J^{2}}{3MV}\right)^{2}\right].$ (2)
This relation admits a single physically meaningful real root for $A>0$ provided that
$\frac{4\pi J^{2}}{3MV}<1.$ (3)
A similar algebraic structure appears in the asymptotically AdS rotating black string solutions obtained by Lemos [33, 34], whose thermodynamic variables obey [36] (see Eq. (47) of Supplement)
$3\mathcal{A}^{6}+2048\pi^{2}\mathcal{J}^{2}\mathcal{V}^{2}-1536\pi\mathcal{V}^{3}\mathcal{M}=0,$ (4)
leading to
${\cal A}^{6}=\frac{512\pi{\cal V}^{2}}{3}\left(3{\cal M}{\cal V}-4\pi{\cal J}^{2}\right).$ (5)
Here again a single positive real solution for $\cal A$} exists only if the inequality (3) is satisfied, demonstrating that the same type of bound governs the existence of the horizon. The similarity of the algebraic constraints highlights a universal feature independently of horizon topology.
[…]
${\cal R}\geq\left[1-\left(\frac{4\pi J^{2}}{3MV}\right)^{2}\right]^{-1/6}.$ (6)
In the limit $J\to 0$, the Kerr-AdS area reduces to the Schwarzschild-AdS area and the RII (${\cal R}\geq 1$) is recovered. The inequality is saturated for the Kerr-AdS black hole. It can be seen from inequality (3) that for the right-hand side of the above inequality we have
${\cal R}_{\rm Kerr-AdS}=\left[1-\left(\frac{4\pi J^{2}}{3MV}\right)^{2}\right]^{-1/6}\geq 1,$ (7)
which verifies the validity of RII in its conventional form in the presence of rotation.
Further validations. For KN-AdS black hole, consider the identity (36). Bring everything to one side and choose the sign so that the cubic term is positive:
$f(A)\equiv M^{2}A^{3}+16\pi^{2}Q^{2}J^{2}A-36\pi M^{2}V^{2}+64\pi^{3}J^{4}=0,$ (8)
which is a cubic equation in $A$. The derivative is
$f^{\prime}(A)=3M^{2}A^{2}+16\pi^{2}Q^{2}J^{2}.$ (9)
For physically meaningful parameters ($M\neq 0$, and $Q,J$ real), this derivative is strictly positive for all real $A$. Therefore, $f(A)$ is a strictly increasing function. That means the equation has exactly one real root. Since $f(A)$ is monotonic, whether the unique root lies at $A>0$ depends only on the sign of $f(0)$. At $A=0$,
$f(0)=-36\pi M^{2}V^{2}+64\pi^{3}J^{4}.$ (10)
So, if $f(0)<0$, then as $A\to\infty$, $f(A)\to\infty$, so the crossing point (the root) must occur at some positive $A$. Therfore, there is exactly one physically meaningful real root with $A>0$ if and only if the inequality (3) is satisfied.
On the other side, rewritting the ratio $4\pi J^{2}/(3MV)$ in terms of geometrical values, reads
$\frac{4\pi J^{2}}{3MV}=\frac{a^{2}\left[\left(a^{2}+r_{+}^{2}\right)\left(r_{+}^{2}+\ell^{2}\right)+q^{2}\ell^{2}\right]}{a^{2}q^{2}\ell^{2}+\left(a^{2}+r_{+}^{2}\right)\left[a^{2}\left(\ell^{2}-r_{+}^{2}\right)+2r_{+}^{2}\ell^{2}\right]}.$ (11)
To make the fraction above is less than one, the numerator must be smaller than the denominator; In other words, the difference between the denominator and numerator of the fraction must be negative, i.e.,
$2r_{+}^{2}\left(a^{2}+r_{+}^{2}\right)\left(a^{2}-\ell^{2}\right)<0,$ (12)
which yields
$\lvert a\rvert$$\ell$, (13)
[…]
For charged and rotating black strings, the ratio is given by
$\frac{4\pi J^{2}}{3MV}$ $=$ $-\frac{r_{+}^{2}\sqrt{a^{2}+\ell^{2}}\left(a^{2}+2\ell^{2}\right)\left(q^{2}\ell^{2}-3r_{+}^{4}\right)\left(q^{2}\ell^{2}+r_{+}^{4}\right)}{12\ell^{2}\left[a^{2}\left(q^{4}\ell^{4}-2q^{2}r_{+}^{4}\ell^{2}+5r_{+}^{8}\right)+\ell^{2}\left(q^{2}\ell^{2}+r_{+}^{4}\right)^{2}\right]}\,.$ (14)
Satisfying the inequality (3) requires that
$q<q_{\rm ext}\sqrt{\frac{a^{2}+\ell^{2}}{3a^{2}+\ell^{2}}},$ (15)
where $q_{\rm ext}$ is the critical value of the black string charge parameter [36],
$q_{\rm ext}=\frac{\sqrt{3}r_{+}^{2}}{\ell},$ (16)
at which, the temperature vanishes, the horizon is degenerate and the black string is extremal. Since for nonzero rotation $a\neq 0$, the ratio
$\frac{a^{2}+\ell^{2}}{3a^{2}+\ell^{2}}<1,$ (17)
hence
$q_{\rm ext}\sqrt{\frac{a^{2}+\ell^{2}}{3a^{2}+\ell^{2}}}<q_{\rm ext},$ (18)
which yields
$q<q_{\rm ext}\sqrt{\frac{a^{2}+\ell^{2}}{3a^{2}+\ell^{2}}}<q_{\rm ext},$ (19)
meaning the black string has a horizon. So, the inequality defines a tighter bound on the charge of the charged and rotating AdS black string.
For our next example, we consider the charged, rotating AdS C-metric. We start from the inequality (55)
$36\pi M^{2}V^{2}-M^{2}A^{3}-64\pi^{3}J^{4}\geq 0\,.$ (20)
Assuming the physical condition $M>0$, we divide the inequality by $M^{2}$ and rewrite it as
$A^{3}\leq 36\pi V^{2}\left[1-\left(\frac{4\pi J^{2}}{3MV}\right)^{2}\right].$ (21)
For the inequality to admit a positive solution for $A$, the right-hand side must be non-negative, because the left-hand side $A^{3}$ is non-negative for all real $A\geq 0$. Thus, the necessary and sufficient condition for the existence of a positive area becomes simply $4\pi J^{2}/(3MV)\leq 1$. Note that if the inequality (3) is satisfied the right-hand side of (21) is strictly positive, and the inequality (21) admits a positive interval of allowed areas, and thus a positive horizon area exists. If $4\pi J^{2}/(3MV)=1$, the right-hand side of (21) becomes zero, and the inequality (21) admits only the trivial solution $A=0$, which is not physically acceptable. Thus no physical positive area exists.
We now turn to rotating black holes with pairwise-equal charges in $D=4$ gauged supergravity. These solutions, characterized by two $U(1)$ charges, were first presented in [37] and later examined from a thermodynamic perspective in [5, 38]. By rewritting the ratio $4\pi J^{2}/(3MV)$ in terms of geometrical values,
$\frac{4\pi J^{2}}{3MV}$ $=$ $a^{2}\left\{\ell^{2}\left[a^{2}+r_{+}(2(q_{1}+q_{2})+r_{+})\right]+r_{1}r_{2}\left(a^{2}+r_{1}r_{2}\right)\right\}\times$ (22)
$\left\{a^{2}\left(a^{2}+r_{1}r_{2}\right)\left[r_{1}r_{2}-2r_{+}(q_{1}+q_{2}+r_{+})\right]\right$.
$\left.+\ell^{2}\left\{a^{4}+a^{2}r_{+}\left[4(q_{1}+q_{2})+3r_{+}\right]+2r_{1}r_{2}r_{+}(q_{1}+q_{2}+r_{+})\right\}\right\}^{-1}.$
In writting the above result, we have used Eqs. (S0.Ex5), (57), and (61). The horizon Eq. (58) has also been used to eliminate the mass parameter $m$. One can check inequality (3) directly by requiring that the numerator be smaller than the denominator. Equivalently, this means that the difference between the denominator and the numerator of the fraction must be negative, i.e.,
$r_{+}(q_{1}+q_{2}+r_{+})\left(a^{2}+r_{1}r_{2}\right)\left(a^{2}-\ell^{2}\right)<0,$ (23)
[…]
Finally, we consider the Kerr-Sen-AdS black hole. Sen discovered a charged, rotating black hole solution in the low-energy limit of heterotic string theory, now commonly referred to as the Kerr–Sen black hole [40]. The thermodynamics is introduced in the extended phase space in [44, 45]. It can be seen from Eq. (73) in the Supplement that the identity satisfied by the thermodynamic variables is analogous to that of the KN-AdS black hole, with the black hole mass $M$ and thermodynamic volume $V$ replaced by $\widetilde{M}$} and $\widetilde{V}$}, respectively. This is because both the standard form of the first law of black hole thermodynamics and the Bekenstein–Smarr mass formula should be satisfied simultaneously. Consequently, the inequality takes the following form:
$\frac{4\pi J^{2}}{3{\widetilde{M}}{\widetilde{V}}}<1.$ (24)
Extension to higher dimensions. In high dimensions, obtaining an identity for thermodynamic variables would generically require solving a polynomial of degree greater than four. This prevents representing the inequality in dimensions larger than four. To obtain the $D$-dimensional version of inequality (3), we consider the intermediate RRII introduced in [32] with this requirement that the RII should also be satisfied in $D$ dimensions in the presence of rotation, which yields
[…]
and
[…]
Supplemental Material
Kerr-Newman-AdS black hole. The metric of the KN-AdS black hole expressed in the Boyer–Lindquist-like coordinates take the form [27]
$ds^{2}$ $\!\!\!\!\!=\!\!\!\!\!$ $-\frac{\Delta_{r}}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}\left(adt-\frac{r^{2}+a^{2}}{\Xi}d\phi\right)^{2},$ (27)
where
$\rho^{2}$ $=$ $r^{2}+a^{2}\cos^{2}\theta,\quad\Xi=1-\frac{a^{2}}{\ell^{2}},\quad\Delta_{r}=(r^{2}+a^{2})\left(1+\frac{r^{2}}{\ell^{2}}\right)-2mr+q^{2},$
$\Delta_{\theta}$ $=$ $1-\frac{a^{2}}{\ell^{2}}\cos^{2}\theta,\quad A=-\frac{qr}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right),$ (28)
and $m$, $q$, and $a$ character the mass $M$, charge $Q$, and angular momentum $J$ of the KN–AdS black hole respectively,
$M=\frac{m}{\Xi^{2}},\quad Q=\frac{q}{\Xi},\quad J=aM=\frac{am}{\Xi^{2}}.$ (29)
Note that requiring a non-singular metric means that $\Xi>0$ which limits $\lvert a\rvert$$\ell$. From Eq. (S0.Ex3), we can fix the horizon radius $r_{+}$ as the largest root of $\Delta_{r=r_{+}}=0$ and simultaneously obtain the mass of the KN–AdS black hole as
$M=\frac{(r_{+}^{2}+a^{2})(r_{+}^{2}+\ell^{2})+q^{2}\ell^{2}}{2r_{+}\ell^{2}\Xi^{2}}.$ (30)
Moreover, the Bekenstein-Hawking entropy $S$ of the KN–AdS black hole can be obtained as
$S=\frac{\pi(r_{+}^{2}+a^{2})}{\Xi}.$ (31)
The extended thermodynamics of KN-AdS black holes was studied in [28, 29]. Solving $r_{+}$ from $S$, substituting it into Eq. (30), and using Eq. (29) and
$P=-\frac{\Lambda}{8\pi}=\frac{3}{8\pi\ell^{2}},$ (32)
we can reexpress the mass of the KN–AdS black hole in terms of the thermodynamic variables, $S$, $P$, $J$, and $Q$, as
$M=\frac{1}{2}\sqrt{\frac{S}{\pi}\left[\left(1+\frac{\pi Q^{2}}{S}+\frac{8PS}{3}\right)^{2}+\frac{4\pi^{2}J^{2}}{S^{2}}\left(1+\frac{8PS}{3}\right)\right]}.$ (33)
From the first law of thermodynamics for the KN-AdS black hole in the extended phase space,
$dM=TdS+VdP+\Omega dJ+\Phi dQ,$ (34)
where $T$, $V$, $\Omega$, and $\Phi$ are the Hawking temperature, thermodynamic volume, angular velocity, and electric potential respectively, we clearly observe that the mass $M$ of the KN-AdS black hole should be essentially identified as enthalpy rather than internal energy. Then, it is straightforward to have
$V$ $=$ $\left(\frac{\partial M}{\partial P}\right)_{S,J,Q}=\frac{2S^{2}}{3\pi M}\left(1+\frac{\pi Q^{2}}{S}+\frac{8PS}{3}+\frac{2\pi^{2}J^{2}}{S^{2}}\right).$ (35)
Solving the equation (35) for pressure and substituting it into the mass formula (33) and using the Bekenstein–Hawking formula $S=A/4$, we can eliminate the pressure and reexpress the mass $M$ in terms of $(A,V,J,Q)$, which leads to the following identity [32]
$36\pi M^{2}V^{2}-M^{2}A^{3}-64\pi^{3}J^{4}=16\pi^{2}Q^{2}J^{2}A.$ (36)
By setting $Q=0$ we find the corresponding identity for Kerr-AdS black hole as
$36\pi M^{2}V^{2}-M^{2}A^{3}-64\pi^{3}J^{4}=0.$ (37)
Asymptotically AdS charged and rotating black string. The metric of asymptotically AdS charged and rotating black string corresponding to cylindrical or toroidal horizons is given by [33, 34]
$ds^{2}=-f(r)\left(\Xi dt-ad\phi\right)^{2}+\frac{1}{f(r)}dr^{2}+\frac{r^{2}}{\ell^{4}}\left(adt-\Xi\ell^{2}d\phi\right)^{2}+\frac{r^{2}}{\ell^{2}}dz^{2},$ (38)
where
$\Xi=\sqrt{1+\frac{a^{2}}{\ell^{2}}},\quad f(r)=\frac{r^{2}}{\ell^{2}}-\frac{m}{r}+\frac{q^{2}}{r^{2}},\quad A=-\frac{q\Xi}{r}\left(dt-\frac{a}{\Xi}d\phi\right).$ (39)
Here, the integration constants $m$ and $q$ are proportional to the mass and charge of the black string, respectively. The constants $a$ and $\ell$, both having dimensions of length, can be interpreted as the rotation parameter and AdS radius, respectively. The relevant thermodynamic potentials per unit length of a stationary black string horizon with cylindrical topology are given by [35]
${\cal M}=\frac{1}{16\pi\ell}\left(3\Xi^{2}-1\right)m,\quad{\cal Q}=\frac{q\Xi}{4\pi\ell},\quad{\cal J}=\frac{3}{16\pi\ell}\Xi am,$ (40)
[…]
${\cal M}=\frac{\left(3\Xi^{2}-1\right)\left(r_{+}^{4}+q^{2}\ell^{2}\right)}{16\pi r_{+}\ell^{3}}.$ (41)
Furthermore, the entropy of black string per unit length of horizon reads
${\cal S}=\frac{r_{+}^{2}\Xi}{4\ell}.$ (42)
The thermodynamic propertises of asymptotically AdS charged and rotating black strings are generalized to the the extended phase space in [36]. Unfortunately, in the case of charged solutions finding an identity among the thermodynamic variables $(\cal M,\cal A,\cal V,\cal J,\cal Q)$ require solving a polynomial of high degree. But, in the case of uncharged solutions, solving $r_{+}$ from $S$, substituting it into Eq. (41), and using Eqs. (40) (by setting $q=0$) and (32) we can reexpress the mass of the black string in terms of the thermodynamic variables, $\cal S$}, $P$ and $\cal J$} as [36],
$\mathcal{M}=\frac{4\sqrt{P}\left(\pi^{3}\mathcal{J}^{4}+\pi^{3/2}\mathcal{J}^{2}\sqrt{\pi^{3}\mathcal{J}^{4}+54P\mathcal{S}^{6}}+18P\mathcal{S}^{6}\right)}{\sqrt{3}\sqrt[4]{\pi}\left(\pi^{3/2}\mathcal{J}^{2}+\sqrt{\pi^{3}\mathcal{J}^{4}+54P\mathcal{S}^{6}}\right)^{3/2}}.$ (43)
From the first law of black hole thermodynamics, is straightforward to obtain the thermodynamic volume per unit length of black string horizon as
$\mathcal{V}=\frac{\sqrt{\pi^{3/2}\mathcal{J}^{2}+\sqrt{\pi^{3}\mathcal{J}^{4}+54P\mathcal{S}^{6}}}}{\sqrt{3}\sqrt[4]{\pi}\sqrt{P}}.$ (44)
Solving the bove equation for pressure leads to
$P=\frac{2\pi\mathcal{J}^{2}}{3\mathcal{V}^{2}}+\frac{6\mathcal{S}^{6}}{\pi\mathcal{V}^{4}}.$ (45)
By plugging the above equation into the mass (43), one can eliminate the pressure and rewrite the mass in terms of $(\cal S,\cal V,\cal J)$
$\mathcal{M}=\frac{4\pi\mathcal{J}^{2}}{3\mathcal{V}}+\frac{8\mathcal{S}^{6}}{\pi\mathcal{V}^{3}}.$ (46)
Using the Bekenstein-Hawking formula, it is straightforward to demonstrate that in terms of thermodynamic variables $(\cal M,\cal A,\cal V,\cal J)$, the following identity holds for asymptototicaly AdS and uncharged rotating black strings [36]
$3\mathcal{A}^{6}+2048\pi^{2}\mathcal{J}^{2}\mathcal{V}^{2}-1536\pi\mathcal{V}^{3}\mathcal{M}=0.$ (47)
Charged and Rotating AdS C-metric. The metric of the charged and rotating AdS C-metric is given by
$ds^{2}=\frac{1}{H^{2}}\left\{-\frac{f}{\Sigma}\left[\frac{dt}{\alpha}-a\sin^{2}\theta\frac{d\phi}{K}\right]^{2}+\frac{\Sigma}{f}dr^{2}+\frac{\Sigma r^{2}}{h}d\theta^{2}+\frac{h\sin^{2}\theta}{\Sigma r^{2}}\left[\frac{adt}{\alpha}-(r^{2}+a^{2})\frac{d\phi}{K}\right]^{2}\right\},$ (48)
where
$f=(1-A^{2}r^{2})\left[1-\frac{2m}{r}+\frac{a^{2}+q^{2}}{r^{2}}+\frac{r^{2}+a^{2}}{\ell^{2}}\right],\quad\Sigma=1+\frac{a^{2}}{r^{2}}\cos^{2}\theta,\quad H=1+Ar\cos\theta,$
$h=1+2mA\cos\theta+\left[A^{2}(a^{2}+q^{2})-\frac{a^{2}}{\ell^{2}}\right]\cos^{2}\theta,\quad\mathcal{A}=-\frac{q}{\Sigma r}\left[\frac{dt}{\alpha}-a\sin^{2}\theta\frac{d\phi}{K}\right]\,.$ (49)
The parameter $A$ is associated with the acceleration of the spacetime. The event horizon is given by the largest root of $f(r_{+})=0$. In the slow-acceleration regime $A\ell<1$, which is the case of interest here, no acceleration horizon appears.
The extended thermodynamics of the charged and rotating AdS C-metric was analyzed in detail in [39], and we have also relied on results from [15]. For further explanation and derivations, we refer the reader to those works.
In the following, we will review the detail of derivation for the identity which holds between thermodynamic variables for charged and rotating AdS C-metric that presented in Ref. [32]. While Eq. (11) of [39] provides the thermodynamic potentials in terms of the metric parameters, in our analysis we found it more practical to use the relations (10) and (11) presented in [15]. The relevant expressions are as follows:
$M^{2}$ $=\frac{\Delta_{\rm C}S}{4\pi}\left[\left(1+\frac{\pi Q^{2}}{\Delta_{\rm C}S}+\frac{8PS}{3\Delta_{\rm C}}\right)^{2}+\left(1+\frac{8PS}{3\Delta_{\rm C}}\right)\left(\frac{4\pi^{2}J^{2}}{\Delta_{\rm C}^{2}S^{2}}-\frac{3C^{2}\Delta_{\rm C}}{2PS}\right)\right]\,$, (50)
$V$ $=\frac{2S^{2}}{3\pi M}\left[\left(1+\frac{\pi Q^{2}}{\Delta_{\rm C}S}+\frac{8PS}{3\Delta_{\rm C}}\right)+\frac{2\pi^{2}J^{2}}{\Delta_{\rm C}^{2}S^{2}}+\frac{9C^{2}\Delta_{\rm C}^{2}}{32P^{2}S^{2}}\right]\,.$ (51)
Here, $M$ denotes the mass, $V$ the thermodynamic volume, $Q$ the electric charge, $J$ the angular momentum and $S$ the entropy of horizon. The remaining $C$ and $\Delta_{\rm C}$ correspond respectively to the average and differential conical deficits of the spacetime, respectively (see Eq. (9) of [15]). An important point for our purposes in the main text is that the parameters satisfy $C\geq 0$ and $\Delta_{\rm C}\geq 0$. We also employed the shorthand notation
$x\equiv\frac{8PS}{\Delta_{\rm C}}\,$, (52)
which is introduced in (13) of [15] and used in (17) of [15]. using the Christodoulou–Ruffini mass formula from [15], specifically equation (17) in that work, which gives:
$\frac{4\pi M^{2}}{S}\leq\left(\frac{3\pi MV}{2S^{2}}-\frac{2C^{2}}{x^{2}}\right)^{2}-4\left(\frac{\pi J}{S}\right)^{4}\,$, (53)
while the combination of (11) and (13) of [15] gives
$\frac{3\pi MV}{2S^{2}}-\frac{2C^{2}}{x^{2}}>0\,.$ (54)
Combining these relations and replacing $S=A/4$ we obtain [32]
$36\pi M^{2}V^{2}-M^{2}A^{3}-64\pi^{3}J^{4}\geq 0\,.$ (55)
Pairwise-equal four-charge AdS black hole in gauged supergravity. These black holes were first constructed in [37], and the explicit form of the metric can be found in that reference. The extended thermodynamic properties were examined in [5, 38]; here, we provide the corresponding thermodynamic quantities.
$M=\frac{m+q_{1}+q_{2}}{\Xi^{2}}\,,\quad Q_{1}=Q_{2}=\frac{\sqrt{q_{1}(q_{1}+m)}}{2\Xi}\,,\quad Q_{3}=Q_{4}=\frac{\sqrt{q_{2}(q_{2}+m)}}{2\Xi}\,$,
$J=\frac{a(m+q_{1}+q_{2})}{\Xi^{2}}\,.$ (56)
Here,
$\Xi=1-a^{2}g^{2}\,,\quad g\equiv\frac{1}{\ell}\,$, (57)
where, $\ell$ is the cosmological length scale. A physically meaningful solution requires that the charges $Q_{1}$ and $Q_{3}$ take real values, and also that $\Xi>0$; in other words, $\lvert a\rvert$$\ell$. The horizon radius $r_{+}$ is determined as the largest root to the following,
$\Delta_{r}=r^{2}+a^{2}-2mr+g^{2}r_{1}r_{2}(r_{2}r_{2}+a^{2})\,.$ (58)
Furthermore, the Bekenstein–Hawking entropy $S$ can be expressed as
$S=\frac{\pi(r_{1}r_{2}+a^{2})}{\Xi}\,$, (59)
where
$r_{1}=r_{+}+2q_{1}\,,\quad r_{2}=r_{+}+2q_{2}\,.$ (60)
From the first law of black hole thermodynamics, the thermodynamic volume can be expressed as [32]
$V$ $=$ $\frac{2\pi}{3\Xi^{2}r_{+}}\left\{2r_{+}\Xi(r_{+}+q_{1}+q_{2})(a^{2}+r_{1}r_{2})\right$. (61)
$\left.+a^{2}\left[2r_{+}(q_{1}+q_{2})+r_{+}^{2}+g^{2}r_{1}^{2}r_{2}^{2}+a^{2}(1+g^{2}r_{1}r_{2})\right]\right\}\,.$
Kerr-Sen-AdS black hole. In the usual Boyer-Lindquist coordinate, the Kerr-Sen-AdS metric reads [44]
$ds^{2}$ $\!\!\!\!\!=\!\!\!\!\!$ $-\frac{\Delta_{r}}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right)^{2}+\frac{\rho^{2}}{\Delta_{r}}dr^{2}+\frac{\rho^{2}}{\Delta_{\theta}}d\theta^{2}+\frac{\sin^{2}\theta\Delta_{\theta}}{\rho^{2}}\left(adt-\frac{r^{2}+2br+a^{2}}{\Xi}d\phi\right)^{2},$
where
$\rho^{2}$ $=$ $r^{2}+2br+a^{2}\cos^{2}\theta,\quad\Delta_{r}=(r^{2}+2br+a^{2})\left(1+\frac{r^{2}+2br}{\ell^{2}}\right)-2mr,$
$\Xi$ $=$ $1-\frac{a^{2}}{\ell^{2}},\quad\Delta_{\theta}=1-\frac{a^{2}}{\ell^{2}}\cos^{2}\theta,\quad A=-\frac{qr}{\rho^{2}}\left(dt-\frac{a\sin^{2}\theta}{\Xi}d\phi\right).$ (63)
The parameter $b$ denotes the dilatonic charge of the black holes and is defined by
$b=\frac{q^{2}}{2m},$ (64)
where $q$ and $m$ represent the electric charge and mass of the black holes, respectively. In the limit $\ell\to\infty$, the metric reduces to the conventional Kerr-Sen black holes. For the non-rotating case ($a=0$), it further simplifies to the Gibbons-Maeda-Garfinkle-Horowitz-Strominger (GMGHS) solution. Gibbons and Maeda [41] originally derived the black hole and black brane solutions for the dilaton field, while Garfinkle, Horowitz, and Strominger [42] extended this to include charged configurations. It is worth noting that four-dimensional Kerr-Sen-(A)dS black holes have been studied from multiple perspectives, including their shadows [43] and phase-space thermodynamics in the extended phase space [45].
We are now in a position to study the thermodynamics of the Kerr-Sen-AdS black hole. Within the framework of the extended phase space, the thermodynamic quantities corresponding to the solution (S0.Ex7) can be computed using standard methods and are given by [44]:
$M$ $=$ $\frac{m}{\Xi^{2}},\quad Q=\frac{q}{\Xi},\quad J=\frac{am}{\Xi^{2}},\quad T=\frac{(r_{+}+b)\left(2r_{+}^{2}+4br_{+}+\ell^{2}+a^{2}\right)-m\ell^{2}}{2\pi\left(r_{+}^{2}+2br_{+}+a^{2}\right)\ell^{2}},$
$S$ $=$ $\frac{\pi\left(r_{+}^{2}+2br_{+}+a^{2}\right)}{\Xi},\quad\Omega=\frac{a\Xi}{\left(r_{+}^{2}+2br_{+}+a^{2}\right)},\quad\Phi=\frac{qr_{+}}{\left(r_{+}^{2}+2br_{+}+a^{2}\right)},$ (65)
where the mass term $m$ can be derived from the condition $\Delta_{r}(r=r_{+})=0$ with $r_{+}$ being the radius of the event horizon.
It is straightforward to verify that the thermodynamic quantities given in Eq. (S0.Ex9) satisfy the Bekenstein–Smarr mass relation
$M=2TS+2\Omega J+\Phi Q-2PV,$ (66)
where the thermodynamic volume $V$, defined as [44]
$V=\frac{4\pi}{3\Xi}\left(r_{+}+b\right)\left(r_{+}^{2}+2br_{+}+a^{2}\right)$ (67)
conjugate to the pressure (32). Unfortunately, the first law reduces merely to a differential identity,
$dM=TdS+\Omega dJ+\Phi dQ-VdP+Jd\Xi/(2a).$ (68)
This occurs because the thermodynamic quantities have been defined in a rotating frame at infinity rather than in the rest frame at infinity.
The Kerr-Sen-AdS solution can be transformed into the rest frame at infinity by performing the simple coordinate transformation $\phi=\phi-at/\ell^{2}$. After a straightforward but tedious computation of the thermodynamic quantities in this frame, one finds that only the mass, the angular velocity, and the thermodynamic volume differ from those given in Eq. (S0.Ex9). These quantities are given by [44]
$\widetilde{M}=M+\frac{a}{\ell^{2}}J,\qquad{\widetilde{\Omega}}=\Omega+\frac{a}{\ell^{2}},\qquad{\widetilde{V}}=V+\frac{4\pi}{3}aJ.$ (69)
In this frame, the thermodynamic quantities consistently satisfy both the standard form of the first law of black hole thermodynamics and the Bekenstein–Smarr mass formula simultaneously,
$d\widetilde{M}$ $=$ $TdS+{\widetilde{\Omega}}dJ+\Phi dQ-{\widetilde{V}}dP,$
$\widetilde{M}$ $=$ $2TS+2{\widetilde{\Omega}}J+\Phi Q-2P{\widetilde{V}}$. (70)
It is straightforward to verify that both the differential and integral mass relations above can be derived from the following Christodoulou–Ruffini–type squared-mass formula [44]:
${\widetilde{M}}^{2}=\left(1+\frac{8PS}{3}\right)\left[\left(1+\frac{8PS}{3}\right)\frac{S}{4\pi}+\frac{\pi J^{2}}{S}+\frac{Q^{2}}{2}\right].$ (71)
From the standard first law of thermodynamics in the extended phase space, given in the first line of Eq. (S0.Ex10), we can obtain the thermodynamic volume of the Kerr-Sen-AdS black hole in terms of thermodynamics potentials as
${\widetilde{V}}=\left(\frac{\partial\widetilde{M}}{\partial P}\right)_{S,J,Q}=\frac{4\sqrt{S}\left\{6\pi^{2}J^{2}+S\left[S(8PS+3)+3\pi Q^{2}\right]\right\}}{3\sqrt{\pi}\sqrt{8PS+3}\sqrt{12\pi^{2}J^{2}+S\left[S(8PS+3)+6\pi Q^{2}\right]}}.$ (72)
Solving the equation (72) for pressure and substituting it into the mass formula (71) and using the Bekenstein–Hawking formula $S=A/4$, one can eliminate the pressure and reexpress the mass $\widetilde{M}$} in terms of $(A,{\widetilde{V}},J,Q)$, which leads to the following identity [32]
$36\pi{\widetilde{M}}^{2}{\widetilde{V}}^{2}-{\widetilde{M}}^{2}A^{3}-64\pi^{3}J^{4}=16\pi^{2}Q^{2}J^{2}A.$ (73)
By setting $Q=0$ we find the corresponding identity for uncharged Kerr-Sen-AdS black hole as
$36\pi{\widetilde{M}}^{2}{\widetilde{V}}^{2}-{\widetilde{M}}^{2}A^{3}-64\pi^{3}J^{4}=0.$ (74)